OpenStudy (anonymous):

What is the center of the circle that has a diameter whose endpoints are (9, 7) and (-3, -5)?

OpenStudy (anonymous):

a. (6, 6) b. (3, 1) c. (3, 4)

OpenStudy (anonymous):

the midpoint of the diameter is the center of your circle. do you know how to find the midpoint?

OpenStudy (anonymous):

no i do not.

OpenStudy (anonymous):

to get the (x0, y0), where x0, y0 is the x, y coordinates respectively, x0 =(x1+x2)/2 y0 = (y1+y2)/2